OpenStudy (anonymous):

Evaluate the function. F(x)= X^3 f(x+(deltax))-f(x)/(deltax) Please show step by step, or atleast how to set it up

OpenStudy (anonymous):

Are you applying the product rule to get the 1st derivative or another one?

OpenStudy (anonymous):

I think this could also be a limits problem. Please confirm the topic of the question

OpenStudy (anonymous):

i think i have to evaluate the f(x)=x^3 into the second one. The answer is 3X^2 +3xdeltax +(deltax)^2 where deltax cannot equal zero. I am just not sure how to set it up

OpenStudy (anonymous):

And it is just evaluating function at the given values. Not a limits problem

OpenStudy (anonymous):

start by computing \[f(x+h)=(x+h)^3\] i.e. multiply out

OpenStudy (anonymous):

you know how to do this?

OpenStudy (anonymous):

when i tried solving it like that i did not get the right answer... I just dont know the right set up

OpenStudy (anonymous):

at the risk of repeating myself, the first step is to cube \[(x+h)^3\] do you know what you get? if not, i can show you, just asking

OpenStudy (anonymous):

is this the first step for the set up? \[f(x)= f(x+Deltax)^3 -x^3/ Deltax\]

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

no one wants to write \(\Delta x\) each time, so lets call it \(h\) which is very common

OpenStudy (anonymous):

ok so f(x+h)^3 -x^3/h

OpenStudy (anonymous):

you job it it to compute \[\frac{f(x+h)-f(x)}{h}\] which has three parts, \(f(x+h), f(x)\) and \(h\)

OpenStudy (anonymous):

after i am done cubing i get: hx^3 + 2hx^2+ 2h^2x^2 + h^2x^3 + h^3x^3/h

OpenStudy (anonymous):

IS that correct or did i mess up cubing?

OpenStudy (anonymous):

i think you might be confused about the notation \[f(x)=x^3\] so for example \[f(2)=2^3\] and \[f(3)=3^3\] and \[f(\pi)=\pi^3\] notice it is not \(f(\pi)=f(\pi)^3\)

OpenStudy (anonymous):

yeah you messed it up

OpenStudy (anonymous):

how tho?

OpenStudy (anonymous):

i am not sure what you did but \[(x+h)^3=x^3+3x^2h+3xh^2+h^3\]

OpenStudy (anonymous):

you can either do \[(x+h)(x+h)(x+h)\] or else just know that \[(a+b)^3=a^3+3a^2b+3ab^2+b^3\]

OpenStudy (anonymous):

ok sorry i got it, i guess the delta x was throwing me off but i rewrote it with h and i got it now, thabk you sooo much

OpenStudy (anonymous):

yw the next job is to subtract \(f(x)=x^3\) that will leave you with \[3x^2h+3xh^2+h^3\] and your final job is to divide every thing by \(h\)

OpenStudy (anonymous):

I got that one too satellite73 LOL...not as fast as you, at first thought I forgot something. Thanks for the calc refresher.

OpenStudy (anonymous):

Do you mind answering another one based on that same concept but with radicals and fractions?

OpenStudy (anonymous):

Satellite - after dividing by h; you have to go back and replace that with delta x, right?